# Inertial antiferromagnetic resonance driven by spin-orbit torques

### v1

@@ IV Frequency and magnitude of resonances
where $\Omega_{1,2}$ are defined in Eqs. (8) and (9). Solving Eq. (16) yields the resonant frequencies, which appear in pairs ±$\pm\omega$. In the absence of spin inertia ($\eta=0$), the resonance condition gives $\omega=\pm\sqrt{\omega_{K}(2\omega_{E}+\omega_{K})}$. In the presence of inertia, the nutational frequency $\omega_{n}$ and the precessional frequency $\omega_{p}$ are given by
@@
In general, resonant peak heights quantify the energy stored in the forced oscillations. As illustrated in Fig. 2, these heights depend on the inertial relaxation time $\eta$ and differ for positive and negative frequencies. Therefore, it is necessary to analyze them in detail. Substituting the resonant frequencies (17) into Eqs. (11)-(12) yields explicit expressions for the peak heights, which are different for each sublattice and each component of magnetization. For the $\omega_{n}$-branch resonance, the peak heights of $x(y)$ components of sublattices $1$ and $2$ are as follows:
$\mathcal{N}_{1,2}^{x(y)}$ $=$ $\rho\frac{\eta}{\sqrt{2}\alpha\kappa_{n}}\sqrt{\frac{\kappa_{n}\!-\!2\eta\omega_{E}}{2\eta\left(\omega_{E}\!+\!\omega_{K}\right)\!+\!\kappa_{n}}}\times$ (19)
$\left|\!\sqrt{\kappa_{n}\!-\!2\eta\omega_{E}}J_{x(y)}\!\mp\!\sqrt{\kappa_{n}\!+\!2\eta\omega_{E}}J_{y(x)}\!\right|,$
@@
$\mathcal{P}_{1,2}^{x(y)}$ $=$ $\rho\frac{\eta}{\sqrt{2}\alpha\kappa_{p}}\sqrt{\frac{\kappa_{p}\!+\!2\eta\omega_{E}}{2\eta\left(\omega_{E}\!+\!\omega_{K}\right)\!-\!\kappa_{p}}}\times$ (20)
$\left|\!\sqrt{\kappa_{p}\!-\!2\eta\omega_{E}}J_{x(y)}\!\pm\!\sqrt{\kappa_{p}\!+\!2\eta\omega_{E}}J_{y(x)}\!\right|,$
The peak heights of the negative ($-\omega_{n}$ and $-\omega_{p}$) branches are obtained by interchanging sublattices $1$ and $2$ in Eqs. (19) and (20), respectively. Throughout the derivation we assume $\alpha^{2}\!\ll\!1$, $\alpha\beta\!\ll\!1$, and $\beta^{2}\!\ll\!1$.
Eqs. (19) and (20) reveal several notable features. (i) The peak heights scale inversely with the damping parameter $\alpha$, a typical characteristic of resonant systems. (ii) The peak heights are highly sensitive to the relative magnitudes of $J_{x}$ and $J_{y}$. For the generic case $J_{x}\neq J_{y}$, as shown in Figs. 3b and 3c, the $x$- and $y$-components exhibit unequal peak heights, indicating elliptical polarization of the resonant modes. (iii) Only when $J_{x}=J_{y}$ do the $x$- and $y$-components exhibit equal peak heights, corresponding to circularly polarized modes. In this case, all four resonances at ±$\pm\omega_{n}$ and ±$\pm\omega_{p}$ exist. This contrasts with the FM case, where only two modes ($-\omega_{n}$ and $\omega_{p}$) appear, as seen in Figs. 7e and 7f. (iv) For specific ratios of $J_{x}$ and $J_{y}$, certain peaks vanish. For example, when $J_{y}/J_{x}=\sqrt{(\kappa_{n}+2\eta\omega_{E})/(\kappa_{n}-2\eta\omega_{E})}$, one finds $\mathcal{N}_{1}^{y}=0$, as indicated by the dashed curve in Fig. 3c. (v) The nutational peaks are generally much smaller than the precessional ones, as evident from comparison between Figs. 3b and 3c. In the limit $\eta=0$, one has $\kappa_{n}=2$ and thus $\mathcal{N}_{1,2}^{x}=\mathcal{N}_{1,2}^{y}=0$, indicating the absence of nutation. Meanwhile, $\kappa_{p}=0$ for $\eta=0$. Then, taking the limit $\eta\rightarrow 0$ in Eq. (20), the precessional peak heights approach
@@
which corresponds to the resonant peak heights of AFM sublattices in the absence of spin inertia.
@@ V Analyses of polarization and handedness
To characterise the resonance polarization we employ two quantities: the ellipticity, which measures the degree of polarization, and the handedness, extracted from the phase difference between $m_{x}$ and $m_{y}$. The ellipticity is defined as the aspect ratio $m_{y}/m_{x}$: it equals $1$ for circular polarization and tends to 00 or $\infty$ for linear polarization; intermediate values correspond to elliptical polarization. For the nutational modes we introduce ellipticity as $e^{n}_{1,2}=\mathcal{N}^{y}_{1,2}/\mathcal{N}^{x}_{1,2}$, and for the precessional modes $e^{p}_{1,2}=\mathcal{P}^{y}_{1,2}/\mathcal{P}^{x}_{1,2}$, where the subscripts $1,2$ label the two sublattices. Using Eqs. (19) and (20) we obtain
$e^{n}_{1,2}$ $=$ $\left|\frac{r_{n}J_{x}\mp J_{y}}{J_{x}\mp r_{n}J_{y}}\right|,$ (22)
$e^{p}_{1,2}$ $=$ $\left|\frac{r_{p}J_{x}\pm J_{y}}{J_{x}\pm r_{p}J_{y}}\right|,$ (23)
@@
In the limit $\eta\to 0$ one recovers $r_{p}=\sqrt{(2\omega_{E}+\omega_{K})/\omega_{K}}$ and $r_{n}=1$. Eqs. (22) and (23) are written for the positive-frequency branches; the corresponding expressions for negative frequencies follow by exchanging the sublattice indices $1\leftrightarrow 2$.
Figure 4: (color online). Ellipticity (upper row) and inter-component phase difference (lower row) as functions of the current ratio $J_{y}/J_{x}$ at positive frequencies for several values of $\eta$. Panels (a)-(d) present the nutational modes of sublattices 1 and 2, while (e)-(h) display the corresponding precessional modes. Solid curves are obtained from Eqs. (22), (23), and (25); symbols denote results calculated without approximations. Parameters are identical to those in Fig. 2.
@@
For the nutational mode, $\mathbf{m}_{1}$ becomes linearly polarized along the $x$-axis ($y$-axis) when $J_{y}/J_{x}=r_{n}$ ($J_{y}/J_{x}=1/r_{n}$), as indicated by the vanishing (diverging) ellipticity $e_{1}^{n}$ in Fig. 4(a). In contrast, Fig. 4(c) shows that $e_{2}^{n}$ remains finite and non-unity at these points, indicating elliptical polarization for $\mathbf{m}_{2}$. Thus, the nutational mode is linearly polarized in one sublattice and elliptically polarized in the other.
A similar feature appears in the precessional mode: at $J_{y}/J_{x}=r_{p}$ ($J_{y}/J_{x}=1/r_{p}$), $\mathbf{m}_{2}$ becomes linearly polarized along the $x$-axis ($y$-axis), while $\mathbf{m}_{1}$ remains elliptically polarized, as seen in Figs. 4(e) and 4(g).
In the limiting case where $J_{x}$ or $J_{y}$ vanishes, all modes are elliptically polarized with ellipticity equal to $r_{n,p}$ or $1/r_{n,p}$. When $J_{x}=J_{y}$, the ellipticity becomes unity, and all modes are circularly polarized. These polarization characteristics are also illustrated schematically in Fig. 5 for representative values of $J_{y}/J_{x}$, in agreement with Fig. 4.
Comparison of the ellipticity curves for different $\eta$ values reveals that increasing spin inertia enhances the deviation of $e^{n}_{1,2}$ and $e^{p}_{1,2}$ from unity, thereby increasing the ellipticity and, thus, the degree of polarization.
Refer to caption Refer to caption
Figure 5: (color online). Polarization and handedness of resonant modes for four branches (±$\pm\omega_{p}$, ±$\pm\omega_{n}$) at various values of $J_{y}/J_{x}$. The inertial relaxation time is set to $\eta=100$ fs, yielding $r_{n}=1.6855$ and $r_{p}=19.4488$ from Eq. (24). In each panel, the horizontal and vertical axes correspond to $m_{x}$ and $m_{y}$ components, respectively; axis ranges are scaled according to the larger amplitude. Curves represent the magnetization orbits, and arrows indicate the sense of rotation.
Besides ellipticity, the handedness – the sense of rotation of $\mathbf{m}_{1,2}$ around the $z$-axis – completes the description of the polarization. Let $\mathbf{m}_{\perp}=m_{k}^{x}\mathbf{e}_{x}+m_{k}^{y}\mathbf{e}_{y}$ denoting the in-plane magnetization. As outlined in Figs. 6a and 6b, a mode is righthanded if $\mathbf{e}_{z}\cdot(\mathbf{m}_{\perp}\times d\mathbf{m}_{\perp}/dt)>0$ or, equivalently, if $\omega\sin(\phi^{x}_{k}-\phi^{y}_{k})>0$ according to Eq. (10). The sign inversion corresponds to the opposite handedness.
@@
$\Delta\phi_{1,2}^{n(p)}\equiv\phi^{x}_{1,2}-\phi^{y}_{1,2}=\tan^{-1}\left(X_{1,2}^{n(p)},Y_{1,2}^{n(p)}\right),$ (25)
@@
$X_{1,2}^{n(p)}$ $=$ ±$\pm\alpha\omega_{n(p)}^{2}\left(J_{y}^{2}-J_{x}^{2}\right),$ (26)
@@
×$\times\left(\sqrt{\kappa_{n}\!+\!2\eta\omega_{E}}J_{x}\!\mp\!\sqrt{\kappa_{n}\!-\!2\eta\omega_{E}}J_{y}\right),$
$Y_{1,2}^{p}\!$ $=$ ±$\!\pm\frac{\kappa_{p}\!+\!2\eta\omega_{E}}{4\eta^{2}}\left(\sqrt{\kappa_{p}\!-\!2\eta\omega_{E}}J_{x}\!\pm\!\sqrt{\kappa_{p}\!+\!2\eta\omega_{E}}J_{y}\right)$ (28)
×$\times\left(\sqrt{\kappa_{p}\!+\!2\eta\omega_{E}}J_{x}\!\pm\!\sqrt{\kappa_{p}\!-\!2\eta\omega_{E}}J_{y}\right).$
where $\kappa_{n,p}$ has been defined in Eq. (18). The phase differences for the $-\omega_{n,p}$ branches follow by exchanging the sublattice indices $1\leftrightarrow 2$ in Eqs. (26)-(28). In deriving Eq. (25) we insert the resonant frequencies for $\omega$ and use the approximations $\alpha^{2}\ll 1$, $\alpha\beta\ll 1$, and $\beta^{2}\ll 1$. For positive resonant frequencies (righthanded SOTs) the mode is righthanded when $0^{\circ}<\Delta\phi_{1,2}^{n(p)}<180^{\circ}$ and lefthanded when $180^{\circ}<\Delta\phi_{1,2}^{n(p)}<360^{\circ}$.
Figure 6: (color online) (a) and (b) Definition of handedness. (c) and (d) Phase diagrams illustrating the polarization of $\mathbf{m}_{1,2}$ as functions of $J_{y}/J_{x}$ and $\eta$ for (c) nutational and (d) precessional resonances. In (c) the upper (lower) boundary is given by $J_{y}/J_{x}=r_{n}$ ($J_{y}/J_{x}=1/r_{n}$); in (d) by $J_{y}/J_{x}=r_{p}$ ($J_{y}/J_{x}=1/r_{p}$). The upper and lower circles in (d) mark $J_{y}/J_{x}=\sqrt{(2\omega_{E}+\omega_{K})/\omega_{K}}$ and $\sqrt{\omega_{K}/(2\omega_{E}+\omega_{K})}$, respectively. The tables above (c) and (d) contain the handedness in different regions for nutational and precessional resonant modes, respectively.
Analogous to the ellipticity, the phase difference $\Delta\phi_{1,2}^{n(p)}$ between the $x$- and $y$-components is governed by the ratio $J_{y}/J_{x}$ (lower panels of Fig. 4). For the nutational branch, the phase difference is $\Delta\phi_{1}^{n}=0^{\circ}$ at $J_{y}/J_{x}=r_{n}$ and $\Delta\phi_{1}^{n}=180^{\circ}$ at $J_{y}/J_{x}=1/r_{n}$ [Fig. 4(b)]. Hence $m_{1}^{x}$ and $m_{1}^{y}$ oscillate in phase at $J_{y}/J_{x}=r_{n}$ or antiphase at $J_{y}/J_{x}=1/r_{n}$, whereas $\mathbf{m}_{1}$ is linearly polarised. In the precessional branch, $\Delta\phi_{2}^{p}=180^{\circ}$ ($\Delta\phi_{2}^{p}=0^{\circ}$) at $J_{y}/J_{x}=r_{p}$ ($J_{y}/J_{x}=1/r_{p}$) [Fig. 4(h)], again indicating that $\mathbf{m}_{2}$ is linear polarized.
@@
For both nutational and precessional modes, the polarization of all resonant modes matches that of the SOTs when the ellipticity of the drive is close to $1$ that is, when $1/r_{n}<J_{y}/J_{x}<r_{n}$ (nutation) or $1/r_{p}<J_{y}/J_{x}<r_{p}$ (precession). This corresponds to Region I in Figs. 6(c) and 6(d). In this regime, both $\mathbf{m}_{1}$ and $\mathbf{m}_{2}$ rotate righthandedly about the $+z$-axis for positive resonant frequencies, and lefthandedly for negative frequencies, consistent with the behavior shown schematically in Fig. 5.
@@
When the ellipticity of the SOTs deviates more strongly from $1$, the polarization behavior becomes sublattice-dependent [Region II in Figs. 6(c) and 6(d)]. Specifically, for nutational modes in Region II ($J_{y}/J_{x}<1/r_{n}$ or $J_{y}/J_{x}>r_{n}$), $\mathbf{m}_{1}$ rotates lefthandedly while $\mathbf{m}_{2}$ maintains righthanded rotation at positive frequency. For precessional modes in Region II ($J_{y}/J_{x}<1/r_{p}$ or $J_{y}/J_{x}>r_{p}$), the handedness is reversed: $\mathbf{m}_{1}$ rotates righthandedly, while $\mathbf{m}_{2}$ rotates lefthandedly for positive frequency.
A comparison between the corresponding panels for positive and negative frequencies in Fig. 5 reveals that flipping the sign of the frequency is equivalent to exchanging $\mathbf{m}_{1}\leftrightarrow\mathbf{m}_{2}$ and simultaneously reversing the rotational sense. This indicates a symmetry of the resonant modes under the combined transformation: sublattice exchange $1\leftrightarrow 2$ and frequency inversion $\omega\leftrightarrow-\omega$.
@@
In FMs, as shown in Appendix B and Fig. 8, the handedness of resonant modes is independent of the ratio $J_{y}/J_{x}$. For positive (negative) frequency, the nutational mode is lefthanded (righthanded), while the precessional mode is righthanded (lefthanded). In contrast, the situation in AFMs is more intricate. As $J_{y}/J_{x}$ crosses the critical values $r_{n(p)}$ or $1/r_{n(p)}$, the handedness of the nutational (precessional) mode is reversed for one of the sublattices. These critical points depend on the inertial relaxation time $\eta$ and are a unique feature of AFMs, absent in the FM case.

### v2

@@ IV Resonant peak and frequency
where $\Omega_{1,2}$ are defined in Eqs. (7) and (8). Solving Eq. (15) yields the resonant frequencies, which appear in pairs $\pm\omega$. In the absence of spin inertia ($\eta=0$), the resonance condition gives $\omega=\pm\sqrt{\omega_{K}(2\omega_{E}+\omega_{K})}$. In the presence of inertia, the nutational frequency $\omega_{n}$ and the precessional frequency $\omega_{p}$ are given by
@@
In general, resonant peak heights quantify the energy stored in the forced oscillations. As illustrated in Fig. 2, these heights depend on the inertial relaxation time $\eta$ and differ for positive and negative frequencies. Therefore, it is necessary to analyze them in detail. Substituting the resonant frequencies (16) into Eqs. (10) and (11) yields explicit expressions for the peak heights, which are different for each sublattice and each component of magnetization. For the $\omega_{n}$-branch resonance, the peak heights of $x(y)$ components of sublattices $1$ and $2$ are as follows:
$\mathcal{N}_{1,2}^{x(y)}$ $=$ $\rho\frac{\eta}{\sqrt{2}\alpha\kappa_{n}}\sqrt{\frac{\kappa_{n}\!-\!2\eta\omega_{E}}{2\eta\left(\omega_{E}\!+\!\omega_{K}\right)\!+\!\kappa_{n}}}$ (18)
$\times\left|\!\sqrt{\kappa_{n}\!-\!2\eta\omega_{E}}J_{x(y)}\!\mp\!\sqrt{\kappa_{n}\!+\!2\eta\omega_{E}}J_{y(x)}\!\right|.$
@@
$\mathcal{P}_{1,2}^{x(y)}$ $=$ $\rho\frac{\eta}{\sqrt{2}\alpha\kappa_{p}}\sqrt{\frac{\kappa_{p}\!+\!2\eta\omega_{E}}{2\eta\left(\omega_{E}\!+\!\omega_{K}\right)\!-\!\kappa_{p}}}$ (19)
$\times\left|\!\sqrt{\kappa_{p}\!-\!2\eta\omega_{E}}J_{x(y)}\!\pm\!\sqrt{\kappa_{p}\!+\!2\eta\omega_{E}}J_{y(x)}\!\right|.$
The peak heights of the negative ($\omega_{n}$ and $\omega_{p}$) branches are obtained by interchanging sublattices $1$ and $2$ in Eqs. (18) and (19), respectively. Throughout the derivation we assume $\alpha^{2}\!\ll\!1$, $\alpha\beta\!\ll\!1$, and $\beta^{2}\!\ll\!1$.
Equations (18) and (19) reveal several notable features. (i) The peak heights are highly sensitive to the relative magnitudes of $J_{x}$ and $J_{y}$. For the generic case $J_{x}\neq J_{y}$, as shown in Figs. 3b and 3c, the $x$ and $y$ components exhibit unequal peak heights, indicating elliptical polarization of the resonant modes. Spin inertia further enhances the ellipticity. (ii) Only when $J_{x}=J_{y}$ do the $x$ and $y$ components exhibit equal peak heights, corresponding to circularly polarized modes. In this case, all four resonances at $\pm\omega_{n}$ and $\pm\omega_{p}$ exist. This contrasts with the FM case, where only two modes ($\omega_{n}$ and $\omega_{p}$) appear, as seen in Figs. 7e and 7f. (iii) The nutational peaks are generally much smaller than the precessional ones, as evident from comparison between Figs. 3b and 3c. In the limit $\eta=0$, one has $\kappa_{n}=2$ and thus $\mathcal{N}_{1,2}^{x}=\mathcal{N}_{1,2}^{y}=0$, indicating the absence of nutation. Meanwhile, $\kappa_{p}=0$ for $\eta=0$. Then taking the limit $\eta\rightarrow 0$ in Eq. (19), the precessional peak heights approach
@@
which corresponds to the resonant peak heights of AFM sublattices in the absence of spin inertia. (iv) For specific ratios of $J_{x}$ and $J_{y}$, certain nutational peak vanishes. For example, when $J_{y}/J_{x}=\sqrt{(\kappa_{n}+2\eta\omega_{E})/(\kappa_{n}-2\eta\omega_{E})}$, one finds $\mathcal{N}_{1}^{y}=0$, as indicated by the brown curve in Fig. 3c. Complete suppression of this nutational peak enables polarization tuning from circular to linear.
@@ V Analyses of polarization and handedness
To characterize the resonance polarization we employ two quantities: the ellipticity, which measures the degree of polarization, and the handedness, extracted from the phase difference between $m_{x}$ and $m_{y}$. The ellipticity is defined as the aspect ratio $m_{y}/m_{x}$: It equals $1$ for circular polarization and tends to 0 or $\infty$ for linear polarization; intermediate values correspond to elliptical polarization. For the nutational modes we introduce ellipticity as $e_{1,2}^{n}=\mathcal{N}_{1,2}^{y}/\mathcal{N}_{1,2}^{x}$, and for the precessional modes $e_{1,2}^{p}=\mathcal{P}_{1,2}^{y}/\mathcal{P}_{1,2}^{x}$, where the subscripts $1,2$ label the two sublattices. Using Eqs. (18) and (19) we obtain
$e_{1,2}^{n}$ $=$ $\left|\frac{r_{n}J_{x}\mp J_{y}}{J_{x}\mp r_{n}J_{y}}\right|,$ (21)
$e_{1,2}^{p}$ $=$ $\left|\frac{r_{p}J_{x}\pm J_{y}}{J_{x}\pm r_{p}J_{y}}\right|,$ (22)
@@
In the limit $\eta\to 0$ one recovers $r_{p}=\sqrt{(2\omega_{E}+\omega_{K})/\omega_{K}}$ and $r_{n}=1$. Equations (21) and (22) are written for the positive-frequency branches; the corresponding expressions for negative frequencies follow by exchanging the sublattice indices $1\leftrightarrow 2$.
Figure 4: (color online). Ellipticity (upper row) and inter-component phase difference (lower row) as functions of the current ratio $J_{y}/J_{x}$ at positive frequencies for several values of $\eta$. Panels (a)-(d) present the nutational modes of sublattices 1 and 2, while (e)-(h) display the corresponding precessional modes. The curves are calculated with Eqs. (21), (22), and (24). Parameters are identical to those in Fig. 2.
@@
For the nutational mode, $\mathbf{m}_{1}$ becomes linearly polarized along the $x$ axis ($y$ axis) when $J_{y}/J_{x}=r_{n}$ ($J_{y}/J_{x}=1/r_{n}$), as indicated by the vanishing (diverging) ellipticity $e_{1}^{n}$ in Fig. 4(a) with red arrows. In contrast, Fig. 4(c) shows that $e_{2}^{n}$ remains finite and nonunity at these points, indicating elliptical polarization for $\mathbf{m}_{2}$. Thus, the nutational mode can be linearly polarized in one sublattice and elliptically polarized in the other.
A similar feature appears in the precessional mode: At $J_{y}/J_{x}=r_{p}$ ($J_{y}/J_{x}=1/r_{p}$), $\mathbf{m}_{2}$ becomes linearly polarized along the $x$-axis ($y$-axis), while $\mathbf{m}_{1}$ remains elliptically polarized, as seen in Figs. 4(e) and 4(g). The linear polarization is indicated with red arrows.
In the limiting case where $J_{x}$ or $J_{y}$ vanishes, all modes are elliptically polarized with ellipticity equal to $r_{n,p}$ or $1/r_{n,p}$. When $J_{x}=J_{y}$, the ellipticity becomes unity, and all modes are circularly polarized. These polarization characteristics are also illustrated schematically in Fig. 5 for representative values of $J_{y}/J_{x}$, in agreement with Fig. 4. Comparison of the ellipticity curves for different $\eta$ values reveals that increasing spin inertia enhances the deviation of $e_{1,2}^{n}$ and $e_{1,2}^{p}$ from unity, thereby increasing the ellipticity and, thus, the degree of polarization.
Figure 5: (color online). Polarization and handedness of resonant modes for four branches ($\pm\omega_{p}$, $\pm\omega_{n}$) at various values of $J_{y}/J_{x}$. The inertial relaxation time is set to $\eta=100$ fs, yielding $r_{n}=1.6855$ and $r_{p}=19.4488$ from Eq. (23). In each panel, the horizontal and vertical axes correspond to $m_{x}$ and $m_{y}$ components, respectively. Curves represent the magnetization orbits, and arrows indicate the sense of rotation. In the panels in columns 1-5, $J_{x}=1$GA//m2 and $J_{y}$ is calculated by the ratio $J_{y}/J_{x}$. In the panels in columns 6-9, $J_{y}=1$GA//m2 and $J_{x}$ is calculated by the ratio.
In addition to ellipticity, the handedness – the sense of rotation of $\mathbf{m}_{1,2}$ around the $z$ axis – completes the description of the polarization. Let $\mathbf{m}_{\perp}=m_{k}^{x}\mathbf{e}_{x}+m_{k}^{y}\mathbf{e}_{y}$ denote the in-plane magnetization. As outlined in Figs. 6a and 6b, a mode is righthanded if $\mathbf{e}_{z}\cdot(\mathbf{m}_{\perp}\times d\mathbf{m}_{\perp}/dt)>0$ or, equivalently, if $\omega\sin(\phi_{k}^{x}-\phi_{k}^{y})>0$ according to Eq. (9). The sign inversion corresponds to the opposite handedness.
@@
$\Delta\phi_{1,2}^{n(p)}=\tan^{-1}\left[X_{1,2}^{n(p)},Y_{1,2}^{n(p)}\right],$ (24)
@@
$X_{1,2}^{n(p)}$ $=$ $\pm\alpha\omega_{n(p)}^{2}\left(J_{y}^{2}-J_{x}^{2}\right),$ (25)
@@
$\times\left(\sqrt{\kappa_{n}\!+\!2\eta\omega_{E}}J_{x}\!\mp\!\sqrt{\kappa_{n}\!-\!2\eta\omega_{E}}J_{y}\right),$
$Y_{1,2}^{p}\!$ $=$ $\!\pm\frac{\kappa_{p}\!+\!2\eta\omega_{E}}{4\eta^{2}}\left(\sqrt{\kappa_{p}\!-\!2\eta\omega_{E}}J_{x}\!\pm\!\sqrt{\kappa_{p}\!+\!2\eta\omega_{E}}J_{y}\right)$ (27)
$\times\left(\sqrt{\kappa_{p}\!+\!2\eta\omega_{E}}J_{x}\!\pm\!\sqrt{\kappa_{p}\!-\!2\eta\omega_{E}}J_{y}\right).$
where $\kappa_{n,p}$ has been defined in Eq. (17). The phase differences for the $\omega_{n,p}$ branches follow by exchanging the sublattice indices $1\leftrightarrow 2$ in Eqs. (25)-(27). In deriving Eq. (24) we insert the resonant frequencies for $\omega$ and use the approximations $\alpha^{2}\ll 1$, $\alpha\beta\ll 1$, and $\beta^{2}\ll 1$. For positive resonant frequencies (righthanded SOTs) the mode is righthanded when $0^{\circ}<\Delta\phi_{1,2}^{n(p)}<180^{\circ}$ and lefthanded when $180^{\circ}<\Delta\phi_{1,2}^{n(p)}<360^{\circ}$.
Figure 6: (color online) (a) and (b) Definition of handedness. Phase diagrams illustrating the polarization of $\mathbf{m}_{1,2}$ as functions of $J_{y}/J_{x}$ and $\eta$ for (c) nutational and (d) precessional resonances. In (c) the upper (lower) boundary is given by $J_{y}/J_{x}=r_{n}$ ($J_{y}/J_{x}=1/r_{n}$); in (d) by $J_{y}/J_{x}=r_{p}$ ($J_{y}/J_{x}=1/r_{p}$). The upper and lower circles in (d) mark $J_{y}/J_{x}=\sqrt{(2\omega_{E}+\omega_{K})/\omega_{K}}$ and $\sqrt{\omega_{K}/(2\omega_{E}+\omega_{K})}$, respectively. The tables above (c) and (d) contain the handedness in different regions for nutational and precessional resonant modes, respectively.
Analogous to the ellipticity, the phase difference $\Delta\phi_{1,2}^{n(p)}$ between the $x$- and $y$-components is governed by the ratio $J_{y}/J_{x}$ (lower panels of Fig. 4). For the nutational branch, the phase difference is $\Delta\phi_{1}^{n}=0^{\circ}$ at $J_{y}/J_{x}=r_{n}$ and $\Delta\phi_{1}^{n}=180^{\circ}$ at $J_{y}/J_{x}=1/r_{n}$ [Fig. 4(b)]. Hence $m_{1}^{x}$ and $m_{1}^{y}$ oscillate in phase at $J_{y}/J_{x}=r_{n}$ or antiphase at $J_{y}/J_{x}=1/r_{n}$, whereas $\mathbf{m}_{1}$ is linearly polarised. In the precessional branch, $\Delta\phi_{2}^{p}=180^{\circ}$ ($\Delta\phi_{2}^{p}=0^{\circ}$) at $J_{y}/J_{x}=r_{p}$ ($J_{y}/J_{x}=1/r_{p}$) [Fig. 4(h)], again indicating that $\mathbf{m}_{2}$ is linear polarized. The nutational mode $\mathbf{m}_{2}$ and precessional mode $\mathbf{m}_{1}$ exhibit the opposite trends with increasing the inertial relaxation time $\eta$, i.e. the phase difference $\Delta\phi_{2}^{n}$ increases with $\eta,$whereas $\Delta\phi_{1}^{p}$ decreases.
@@
For both nutational and precessional modes, the polarization of all resonant modes matches that of the SOTs when the ellipticity of the drive is close to $1$ that is, when $1/r_{n}<J_{y}/J_{x}<r_{n}$ (nutation) or $1/r_{p}<J_{y}/J_{x}<r_{p}$ (precession). This corresponds to Region I in Figs. 6(c) and 6(d). In this regime, both $\mathbf{m}_{1}$ and $\mathbf{m}_{2}$ rotate righthandedly about the +z+z axis for positive resonant frequencies, and lefthandedly for negative frequencies, consistent with the behavior shown schematically in Fig. 5.
@@
When the ellipticity of the SOTs deviates more strongly from $1$, the polarization behavior becomes sublattice dependent [Region II in Figs. 6(c) and 6(d)]. Specifically, for nutational modes in Region II ($J_{y}/J_{x}<1/r_{n}$ or $J_{y}/J_{x}>r_{n}$), $\mathbf{m}_{1}$ rotates lefthandedly while $\mathbf{m}_{2}$ maintains righthanded rotation at positive frequency. For precessional modes in Region II ($J_{y}/J_{x}<1/r_{p}$ or $J_{y}/J_{x}>r_{p}$), the handedness is reversed: $\mathbf{m}_{1}$ rotates righthandedly, while $\mathbf{m}_{2}$ rotates lefthandedly for positive frequency. We emphasize that the boundaries of the polarization regions depend sensitively on the inertial relaxation time $\eta$, enabling its experimental determination. By tuning the ratio of two orthogonally applied currents and measuring the resulting mode polarizations, $\eta$ can be extracted by fitting the data with Eqs. (21-23).
From Figs. 6(c) and 6(d) we infer that Region I of the nutational mode is much smaller than that of the precessional mode because $r_{n}\ll r_{p}$. This is consistent with Eq. (23), which for $\omega_{K}/\omega_{E}\ll 1$ gives
$r_{n}$ $=\sqrt{1+2\eta\omega_{E}}+O\left[\omega_{K}\right],$ (28)
$r_{p}$ $=\sqrt{\frac{2\omega_{E}}{\omega_{K}}}\sqrt{1+2\eta\omega_{E}}+O\left[\omega_{K}^{1/2}\right].$
For typical AFM parameters with $\omega_{K}\ll\omega_{E}$, it follows that $r_{n}\ll r_{p}$. Consequently, the handedness of the nutational mode reverses at a much smaller current difference than that of the precessional mode.
@@
In FMs, as shown in Appendix C and Fig. 8, the handedness of resonant modes is independent of the ratio $J_{y}/J_{x}$. For positive (negative) frequency, the nutational mode is lefthanded (righthanded), while the precessional mode is righthanded (lefthanded). In contrast, the situation in AFMs is more intricate. As $J_{y}/J_{x}$ crosses the critical values $r_{n(p)}$ or $1/r_{n(p)}$, the handedness of the nutational (precessional) mode is reversed for one of the sublattices. These critical points depend on the inertial relaxation time $\eta$ and are a unique feature of AFMs, absent in the FM case.